# ”They’ve Stolen My GPL-Licensed Model!”: Toward Standardized and Transparent Model Licensing

### v1

@@ 3.3. Compliance Analysis
Appendix C presents the logic code for rights granting analysis, and the full list of reported notices, warnings, and errors is shown in Table 2. Due to its staged logic rule reasoning design, MG Analyzer has considerable extensibility, enabling the incorporation of additional licenses and analysis targets, provided that it can reason based on our proposed dependencies information. Beyond compliance analysis, our vision is to promote ML workflow supply chain management and improve FAIRness (Wilkinson et al., 2016) in model publishing. We position our vocabulary and tool as a first step toward Linked Open Model Production Data (Bizer et al., 2008).
@@ 4. MG Licenses and Comparison Results
Table 3. List of MG Analyzer-Supported Licenses & Aggrements (including MG Licenses V1.0) with Their Comparisons in Clarity and Freedom. Grouped by OSS, Free-Content (&Dataset), Model and sorted first by clarity score, then by freedom score.
@@
AGPL-3.0 $\bm{\approx}$${\approx}$${\approx}$${\approx}$${\approx}$${\approx}$ 3.5 4.5
AFL-3.0 $\bm{\approx}$${\approx}$${\approx}$${\approx}$${\approx}$ 3.0 7.5
OSL-3.0 $\bm{\approx}$${\approx}$${\approx}$${\approx}$ 3.0 5
Apache-2.0 $\bm{\approx}$${\approx}$${\approx}$${\approx}$ 2.5 7.5
LGPL-3.0 $\bm{\approx}$${\approx}$${\approx}$${\approx}$ 2.5 7.5
Artistic-2.0 $\bm{\approx}$${\approx}$${\approx}$${\approx}$${\approx}$ 2.5 6.0
GPL-3.0 $\bm{\approx}$${\approx}$${\approx}$${\approx}$${\approx}$${\approx}$ 2.5 4.5
ECL-2.0 $\bm{\approx}$${\approx}$${\approx}$${\approx}$${\approx}$ 2.0 7.5
Unlicensed $\bm{\approx}$${\approx}$${\approx}$ 1.5 10
MIT $\bm{\approx}$${\approx}$${\approx}$${\approx}$ 1.5 8.5
GPL-2.0 $\bm{\approx}$${\approx}$${\approx}$${\approx}$ 1.5 5.5
LGPL-2.1 $\bm{\approx}$${\approx}$${\approx}$${\approx}$ 1.5 5.5
BSD-3-Clause $\bm{\approx}$${\approx}$${\approx}$${\approx}$${\approx}$ 1.0 7.5
BSD-3-Clause-Clear $\bm{\approx}$${\approx}$${\approx}$${\approx}$${\approx}$ 1.0 7.5
BSD-2-Clause $\bm{\approx}$${\approx}$${\approx}$${\approx}$${\approx}$ 1.0 7.5
@@
ODC-By-1.0 $\bm{\approx}$${\approx}$${\approx}$${\approx}$${\approx}$ 3.0 7.5
PDDL-1.0 $\bm{\approx}$ 2.5 10
CC-BY-4.0 $\bm{\approx}$${\approx}$${\approx}$${\approx}$${\approx}$${\approx}$${\approx}$ 2.5 6.0
CC-BY-SA-4.0 $\bm{\approx}$${\approx}$${\approx}$${\approx}$${\approx}$${\approx}$${\approx}$ 2.5 6.0
CC-BY-NC-4.0 $\bm{\approx}$${\approx}$${\approx}$${\approx}$${\approx}$${\approx}$${\approx}$ 2.5 5.0
CC-BY-NC-SA-4.0 $\bm{\approx}$${\approx}$${\approx}$${\approx}$${\approx}$${\approx}$${\approx}$ 2.5 5.0
CC-BY-ND-4.0 $\bm{\approx}$${\approx}$${\approx}$ n/a** n/a n/a n/a 2.5 4.0
CC-BY-NC-ND-4.0 $\bm{\approx}$${\approx}$${\approx}$ n/a** n/a n/a n/a 2.5 3.0
GFDL $\bm{\approx}$${\approx}$${\approx}$${\approx}$${\approx}$ 2.0 3.5
C-UDA $\bm{\approx}$${\approx}$ 1.5 5.5
LGPLLR $\bm{\approx}$${\approx}$${\approx}$${\approx}$${\approx}$${\approx}$ 1.5 4.5
MG0 $\bm{\approx}$${\approx}$${\approx}$ 4.0 8.5
MG-BY $\bm{\approx}$${\approx}$${\approx}$ 4.0 7.5
MG-BY-RAI $\bm{\approx}$${\approx}$${\approx}$ 4.0 6.5
‡ OpenRAIL-M $\bm{\approx}$${\approx}$${\approx}$ 4.0 6.5
MG-BY-NC $\bm{\approx}$${\approx}$${\approx}$${\approx}$ 4.0 6.0
MG-BY-OS $\bm{\approx}$${\approx}$ 4.0 6.0
MG-BY-NC-RAI $\bm{\approx}$${\approx}$${\approx}$${\approx}$ 4.0 5.0
MG-BY-NC-OS $\bm{\approx}$${\approx}$${\approx}$ 4.0 4.5
MG-BY-ND $\bm{\approx}$ n/a n/a n/a n/a 4.0 4.5
MG-BY-NC-ND $\bm{\approx}$ n/a n/a n/a n/a 4.0 3.5
† OPT-175B $\bm{\approx}$${\approx}$${\approx}$ 3.5 5.0
† Llama3 $\bm{\approx}$${\approx}$${\approx}$${\approx}$ 2.5 5.5
† Llama3.1 $\bm{\approx}$${\approx}$${\approx}$${\approx}$ 2.5 5.5
† • AI2-ImpACT-LR $\bm{\approx}$${\approx}$${\approx}$${\approx}$ 2.5 5.5
† • AI2-ImpACT-MR $\bm{\approx}$ n/a n/a $\bm{\approx}$ 2.5 2.0
† • AI2-ImpACT-HR $\bm{\approx}$ n/a n/a n/a n/a n/a n/a 2.5 0
† Gemma $\bm{\approx}$${\approx}$${\approx}$${\approx}$ 2.0 4.5
† Llama2 $\bm{\approx}$${\approx}$${\approx}$${\approx}$ 1.5 5.5
@@
Prefixes: ✓ The license explicitly includes sufficient prefixes that clearly describe scope and conditions of granting rights (e.g., revocable, sublicensable); $\bm{\approx}$ Some important prefixes are indeterminate; ✗ No prefixes are declared.
Rights: ✓ The license explicitly declares whether a patent license or a copyright license is granted; $\bm{\approx}$ Only the granting of a patent license or copyright license is stated; ✗ No explicit grant of either is provided.
Rules: ✓ The license terms cover all actions listed in Table 1; $\bm{\approx}$ Some actions fall outside the definition of this license; ✗ Almost no rules are set forth.
@@
Share: ✓ The license permits the sharing of verbatim copies/derivatives created by you without any restrictions; $\bm{\approx}$ Some restrictions apply to sharing; ✗ Sharing verbatim copies/derivatives is prohibited.
Close: ✓ The license does not require you to disclose the source files of verbatim copies/derivatives created by you; $\bm{\approx}$ Modification statements are required; ✗ You must disclose the source files of your created copies/derivatives.
Non-exclusive: ✓ The license does not restrict you from adding new terms when republishing; $\bm{\approx}$ Certain types of terms are prohibited in republishing; ✗ All republishing must adhere to the original terms and conditions.
Sublicense: ✓ The license explicitly grants sublicensing rights; $\bm{\approx}$ The license prohibits sublicensing but offers automatic licensing instead; ✗ Sublicensing is either prohibited or not explicitly permitted.
Attribute: ✓ The license does not require retaining the original attribution and licenses in redistributed derivatives; $\bm{\approx}$ Attribution or license must be retained; ✗ Redistributed derivatives must retain the attributions and licenses.
Commercial: ✓ The license explicitly grants commercial rights; $\bm{\approx}$ Commercial rights are not explicitly granted but not reserved either, or compromised commercial rights are granted; ✗ Commercial rights are reserved.
Behavioral: ✓ The license does not restrict user behaviors; $\bm{\approx}$ Includes runtime controls (e.g., forced updates); ✗ Certain behaviors involving the licensed materials or derivatives are prohibited (e.g., harming, medical advice).
Clarity/Freedom Score: ✓ $+1.0$, $\bm{\approx}$ $+0.5$, ✗ $+0$, n/a: $+0$ Maximum Clarity Score: 4.0, Maximum Freedom Score: 10.
@@
To reveal the underlying dilemmas in model licensing, we provide comprehensive comparisons of these licenses in Table 3. Based on the terms of these licenses, we evaluated each license’s clarity score and freedom score, each encompassing sub-items as defined in the table. A higher clarity score indicates that the license is more clearly defined in model publishing scenarios, while a higher freedom555 Our freedom score reflects only the amount of restrictions stipulated in a license and should not be confused with the definitions of ”Freedom” in free software (Perens, 1999). score signifies fewer restrictions on republished copies and derivatives. The significant findings are summarized below:
@@
For free-content and dataset licenses, their average clarity score is only slightly better than that of OSS licenses, and they also lack coverage for publishing as a service. However, some CC licenses, such as CC-BY-NC-4.0, offer additional options that prohibit the commercial use of the work, unlike OSS licenses. This may explain why many models are published under these licenses666 There are 9,730 models on HuggingFace licensed under CC-BY-NC-4.0 (more than Llama 2), and 770 of them have garnered 1k+ downloads, please ref to Appendix B., despite the fact that they were not originally drafted for models.
@@
Based on the above findings, we conclude that it is necessary to draft new standardized and flexible licenses for general model publishing. To address this, we collaborated with a law firm to draft a set of model licenses, tentatively referred to as MG Licenses888 See MG License text at this link.. As reflecting in Table 3, our licenses include specific definitions and terms related to ML concepts, offering greater clarity compared to OSS and free-content licenses. Most importantly, following the philosophy of CC licenses, MG Licenses are designed to be flexible and easy to use, offering five options to accommodate various publishing scenarios. The options include: BY (Attribution), NC (Non-Commercial), ND (No Derivatives), RAI (Responsible Use of AI), and OS (Open Source)999 At the time of writing, the Open Source Initiative has only released a draft v. 0.0.9 version of the Open Source AI definition, and our MG licenses with the OS option are not OSI-approved at present.. We drafted nine preset licenses using these options, including MG0, which does not apply any options. The flexibility of our licenses is reflected in table 3, where their freedom scores are evenly distributed between 3.5 and 8.5, indicating that they form a superset of all other model licenses101010 AI2-ImpACT-MR and AI2-ImpACT-HR prohibit sharing copies, we do not consider such restrictions to be common in model publishing, so we have excluded them. For details on how to substitute other licenses with MG licenses, refer to Appendix D.. To promote transparency, we introduced a Model Sheet as an attachment to each MG License, inspired by MDL (Benjamin et al., 2019). This sheet assists model users in understanding the rights and restrictions granted by the license terms and helps model developers identify the most suitable license for their needs (ref Appendix D).
In the next section, we explore which licenses can protect your model from misuse beyond your intended purposes. For example, licensee may close source your GPL-licensed model without violating your license, even if it does contains code disclosure clauses. To do this, we evaluate some of the popular licenses for model publishing, as well as the MG Licenses, using MG Analyzer.
@@ 5. Preliminary License Analysis Results
This section seeks to answer a key question: Should I continue using traditional OSS and free-content licenses to publish my model, and what are the associated risks? To explore this question, we evaluate commonly adopted licenses in the context of model publishing by MG Analyzer to assess whether they are still effective as intended. The example workflow involves combining two models and publishing them as a service, as shown in Figure 4. Here, we consider two scenarios: 1) publishing the combined model as a service; 2) publishing the data generated from the combined model. Models A and B are non-binding works that correspond to the settings in Table 4, with their respective analysis results also presented in the table. Model C is an intermediate work created by combining A and B, and Data D is the generated output from C. Work E involves republishing C as a service with the intent to sell, while Work F involves republishing D as a literary form, also with the intent to sell. To be more convincing, we use real-world models and their respective licenses for demonstration.
@@
C: Unlicense. D: Unlicense. E: W1$\times$2; E2$\times$2; E5$\times$2. F: W1$\times$2 .
@@
First, we evaluate two OSS licenses: AGPL-3.0 and GPL-3.0, which are considered enforceable open-source licenses with copyleft clauses. In setting (i), Work E triggers two Disclose Source Code warnings (code W5, refer to Table 2 for code definitions), and AGPL-3.0 successfully proliferates [^11] While two copyleft conditions are simultaneously triggered here, they can be resolved because GPL-3.0 is compatible with AGPL-3.0. Appendix E visualizes this workflow. to Work C. However, with a small adjustment, we can circumvent these clauses by republishing the generated content rather than providing the model as a service. As demonstrated by Work D, there is no W5 warning, and the content is licensed under the Unlicense. Furthermore, the condition in AGPL-3.0 that triggers the disclose code clauses related to remote access is you modify the Program, which means you can directly republish copies as a service to circumvent this clause. As reflected in setting (ii), there are no more W5 warnings, only two warnings related to the non-standard licensing remain.
@@
Third, we evaluate MG Licenses: MG-BY-OS and MG-BY-NC, which contain open sourcing and non-commercial use clauses, respectively. In setting (iii), our MG-BY-OS license successfully triggers the W5 warning, indicating that Work E must disclose its source code. In setting (iv), the non-commercial use error E5 is reported by the generated Work F. As a model license, we do not enforce licensing on generated content, allowing Work D to be licensed under the Unlicense, albeit with certain restrictions. A summary of the rights granted and restrictions imposed by these licenses can be found in their Model Sheet provided in Appendix D.
It is worth mentioning that all results were obtained with fuzzy form matching enabled, maximizing the detection of potential risks. If these fuzz rules were disabled, fewer issues would be reported. Furthermore, our MG Analyzer is designed to help developers be aware of potential compliance issues in ML projects. Its results should not be considered legal advice or a defense in dispute resolution. Please refer to our disclaimers in Appendix A.
@@ 6. Conclusion
Appendix A Disclaimers
The information in this article is for general informational purposes only and does not constitute legal advice. Views, opinions, and recommendations expressed are solely those of the author(s) and do not represent any organization. Do not rely on this material as a substitute for professional legal advice tailored to your specific circumstances.
Appendix B Count of Licenses in Model Publishing
Table 5 shows the count of licenses used for model publishing listed on HuggingFace. Non-standard model licensing are quite common, as many models are published under OSS licenses and free-content licenses. The MG Analyzer-supported licenses are marked in bold. Although earlier versions (2.0, 3.0, etc.) of CC licenses are marked as not supported, these licenses are similar to the 4.0 version and can be considered as supported.
Table 5. Count of Licenses Used by Models on HuggingFace (Accessed on October 14, 2024). The table includes OSS, Free-content, and Model licenses. Licenses supported by the MG Analyzer are indicated in bold.
Model License Count (Total: 1,050,569)
License Name Count License Name Count
Apache-2.0 152,317 LGPL-3.0 119
MIT 61,974 CDLA-Permissive-2.0 110
Other 33,532 BSD-2-Clause 101
OpenRAIL 32,712 C-UDA 101
CreativeML-OpenRAIL-M 23,653 OSL-3.0 98
CC-BY-NC-4.0 9,730 CC-BY-ND-4.0 90
Llama2 6,971 CC-BY-NC-3.0 83
Llama3 4,970 Ms-PL 67
CC-BY-4.0 4,954 PDDL 62
OpenRAIL++ 4,120 ECL-2.0 59
Gemma 2,973 GFDL 52
CC-BY-NC-SA-4.0 2,738 BSD-3-Clause-Clear 51
AFL-3.0 2,675 CC-BY-NC-ND-3.0 48
Llama3.1 2,323 Zlib 41
CC-BY-SA-4.0 1,974 CC-BY-NC-SA-3.0 39
GPL-3.0 1,968 MPL-2.0 38
CC 1,346 CC-BY-NC-SA-2.0 35
Artistic-2.0 1,030 ODBL 33
BSD-3-Clause 900 DeepFloyd-IF-License 32
BigScience-OpenRAIL-M 861 LGPL 30
BigScience-BLOOM-RAIL-1.0 787 CC-BY-2.5 28
CC-BY-NC-ND-4.0 753 ODC-by 27
BigCode-OpenRAIL-M 686 LGPL-LR 26
WTFPL 662 EPL-2.0 25
Llama3.2 630 ISC 22
AGPL-3.0 595 Apple-ASCL 22
CC0-1.0 508 EUPL-1.1 20
Unlicensed 443 NCSA 16
CC-BY-SA-3.0 356 CDLA-Sharing-1.0 16
GPL 295 Etalab-2.0 16
BSD 241 LGPL-2.1 13
CC-BY-NC-2.0 181 PostgreSQL 12
CC-BY-2.0 153 LPPL-1.3c 11
CC-BY-3.0 143 OFL-1.1 9
GPL-2.0 134 EPL-1.0 9
BSL-1.0 134 CDLA-Permissive-1.0 6
Appendix C Encoded Rules and Reasoning Logic
In this section, we present the encoded ”derivative” rules for the GPL-3.0 and Llama 2 licenses in Turtle format, followed by a snippet of the reasoning logic employed for license determination and rights granting analysis in Notation3.

### v2

@@ 3.3. Compliance Analysis
Appendix B presents the logic code for rights granting analysis, and the full list of reported notices, warnings, and errors is shown in Table 1. Due to its staged logic rule reasoning design, MG Analyzer has considerable extensibility, enabling the incorporation of additional licenses and analysis targets, provided that it can reason based on our proposed dependencies information. Beyond compliance analysis, our vision is to promote ML workflow supply chain management and improve FAIRness (Wilkinson et al., 2016) in model publishing. We position our vocabulary and tool as a first step toward Linked Open Model Production Data (Bizer et al., 2008).
@@ 4. MG Licenses and Comparison Results
Table 2. List of MG Analyzer-Supported Licenses & Agreements (including MG Licenses V1.0) with Their Comparisons in Clarity and Freedom. Grouped by OSS, Free-Content (&Dataset), Model and sorted first by clarity score, then by freedom score.
@@
[…]
AFL-$\bm{\approx} \bm{\approx} \bm{\approx} \bm{\approx} \bm{\approx} 3.0$ 7.5
OSL-$\bm{\approx} \bm{\approx} \bm{\approx} \bm{\approx} 3.0$ 5
[…]
ECL-$\bm{\approx} \bm{\approx} \bm{\approx} \bm{\approx} \bm{\approx} 2.0$ 7.5
[…]
@@
[…]
@@
[…]
@@
[…]
@@
To reveal the underlying dilemmas in model licensing, we provide comprehensive comparisons of these licenses in Table 2. Based on the terms of these licenses, we evaluated each license’s clarity score and freedom score, each encompassing sub-items as defined in the table. A higher clarity score indicates that the license is more clearly defined in model publishing scenarios, while a higher freedom [^9] Our freedom score reflects only the amount of restrictions stipulated in a license and should not be confused with the definitions of ”Freedom” in free software (Perens, 1999). score signifies fewer restrictions on republished copies and derivatives. The significant findings are summarized below:
@@
For free-content and dataset licenses, their average clarity score is only slightly better than that of OSS licenses, and they also lack coverage for publishing as a service. However, some CC licenses, such as CC-BY-NC-4.0, offer additional options that prohibit the commercial use of the work, unlike OSS licenses. This may explain why many models are published under these licenses [^10] There are 9,730 models on HuggingFace licensed under CC-BY-NC-4.0 (more than Llama 2), and 770 of them have garnered 1k+ downloads (Accessed on May 29. 2025)., despite the fact that they were not originally drafted for models.
@@
Based on the above findings, we conclude that it is necessary to draft new standardized and flexible licenses for general model publishing. To address this, we collaborated with a law firm to draft a set of model licenses, tentatively referred to as MG Licenses. As reflecting in Table 2, our licenses include specific definitions and terms related to ML concepts, offering greater clarity compared to OSS and free-content licenses. Most importantly, following the philosophy of CC licenses, MG Licenses are designed to be flexible and easy to use, offering five options to accommodate various publishing scenarios. The options include: BY (Attribution), NC (Non-Commercial), ND (No Derivatives), RAI (Responsible Use of AI), and OS (Open Source). We drafted nine preset licenses using these options, including MG0, which does not apply any options. The flexibility of our licenses is reflected in Table 2, where their freedom scores are evenly distributed between 3.5 and 8.5, indicating that they form a superset of all other model licenses [^12] AI2-ImpACT-MR and AI2-ImpACT-HR prohibit sharing copies, we do not consider such restrictions to be common in model publishing, so we have excluded them. For details on how to substitute other licenses with MG licenses, refer to Appendix C.. To promote transparency, we introduced a Model Sheet as an attachment to each MG License, inspired by MDL (Benjamin et al., 2019). This sheet assists model users in understanding the rights and restrictions granted by the license terms and helps model developers identify the most suitable license for their needs (see Attachment A of MG0-1.0).
In the next section, we examine which licenses protect models from misuse beyond their intended purposes (e.g., allowing closed-sourcing under GPL). We evaluate popular model licenses and the MG Licenses using MG Analyzer.
@@ 5. Preliminary License Analysis Results
This section seeks to answer a key question: Should I continue using traditional OSS and free-content licenses to publish my model, and what are the associated risks? We evaluate commonly adopted licenses using MG Analyzer to assess their effectiveness in this context. The example workflow involves combining two models and publishing them as a service, as shown in Figure 4. Here, we consider two scenarios: 1) publishing the combined model as a service; 2) publishing the data generated from the combined model. Models A and B are non-binding works that correspond to the settings in Table 3, with their respective analysis results also presented in the table. Model C is an intermediate work created by combining A and B, and Data D is the generated output from C. Work E involves republishing C as a service with the intent to sell, while Work F involves republishing D as a literary form, also with the intent to sell. To be more convincing, we use real-world models and their respective licenses for demonstration.
@@
[…]
@@
First, we evaluate two OSS licenses: AGPL-3.0 and GPL-3.0, which are considered enforceable open-source licenses with copyleft clauses. In setting (i), Work E triggers two Disclose Source Code warnings (code W5, refer to Table 1 for code definitions), and AGPL-3.0 successfully proliferates [^13] Although two copyleft conditions are triggered simultaneously, they can be resolved since GPL-3.0 is compatible with AGPL-3.0. Appendix E visualizes this workflow. to Work C. However, with a small adjustment, we can circumvent these clauses by republishing the generated content rather than providing the model as a service. As demonstrated by Work D, there is no W5 warning, and the content is licensed under the Unlicense. Furthermore, the condition in AGPL-3.0 that triggers the disclose code clauses related to remote access is you modify the Program, which means you can directly republish copies as a service to circumvent this clause. As reflected in setting (ii), there are no more W5 warnings, only two warnings related to the non-standard licensing remain.
@@
Third, we evaluate MG Licenses: MG-BY-OS and MG-BY-NC, which contain open sourcing and non-commercial use clauses, respectively. In setting (iii), our MG-BY-OS license successfully triggers the W5 warning, indicating that Work E must disclose its source code. In setting (iv), the non-commercial use error E5 is reported by the generated Work F. As a model license, we do not enforce licensing on generated content, allowing Work D to be licensed under the Unlicense, albeit with certain restrictions. A summary of the rights granted and restrictions imposed by these licenses can be found in their Model Sheet, as previously mentioned.
It is worth mentioning that all results were obtained with fuzzy form matching enabled, maximizing the detection of potential risks. If these fuzz rules were disabled, fewer issues would be reported. Furthermore, our MG Analyzer is designed to help developers be aware of potential compliance issues in ML projects. Its results should not be considered legal advice or a defense in dispute resolution. Please refer to our disclaimers in Appendix A. The most promising application lies in providing license analysis within complex ML workflows. An in-the-wild case can be found in Appendix D.
@@ 6. Conclusion
Acknowledgements.
This research is sponsored by Shanghai Pujiang Programme (Award No. 25PJA029). This research is also supported in part by the National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority.
Appendix A Limitations
The main limitation of MG Analyzer is that it is relied on accurate description of ML workflow production. However, there is currently no standardized framework for capturing and organizing such information, which is typically scattered across Model Cards and Datasheets (Biderman et al., 2022) in unstructured natural language form. Furthermore, even model-sharing platforms like Hugging Face provide user-annotated metadata such as ”trained on” and ”merged from,” but this information is often low quality and incomplete. This limitation hinders the application of our analyzer to large-scale compliance analysis of real-world ML workflows and highlights the need for future efforts to build a comprehensive database that can trace the evolution of ML assets. Another challenge arises from ambiguity in legal interpretations and dispute resolution. Take SmolLM-360M-Instruct in Figure 6 as an example. It is finetuned on data generated by Llama 3.1, which can be interpreted as a form of distillation, potentially rendering it a derivative work of Llama 3.1. However, an opposing perspective argues that since Llama 3.1 was not directly used during training, the resulting model may not be subject to the terms of its license agreement. This legal uncertainty limits the analyzer’s ability to provide definitive compliance assessments in complex or indirect reuse scenarios.
There are also some critical voices regarding the Model Sheet included in our proposed MG Licenses, as noted by several legal experts. While meant to aid understanding, including it in the main license text may create ambiguity and could encourage unauthorized modifications. Currently, MG Licenses are being actively updated based on feedback, with several variants submitted to the OSI for approval. Please see the disclaimers that follow.
Disclaimers: The information in this article is for general informational purposes only and does not constitute legal advice. Views, opinions, and recommendations expressed are solely those of the author(s) and do not represent any organization. Do not rely on this material as a substitute for professional legal advice tailored to your specific circumstances.
Appendix B Encoded Rules and Reasoning Logic
In this section, we present the one encoded ”derivative” rule for the GPL-3.0 license in Turtle format, followed by a snippet of the reasoning logic employed for license determination and rights granting analysis in Notation3.